OpenStudy (anonymous):

The height and width of a rectangular prism are each 2 inches shorter than the length of the prism. The volume of the prism is 40 cubic inches. Approximate the dimensions of the prism to the nearest hundredth.

OpenStudy (anonymous):

h = L - 2 w = L - 2 volume = L*w*h, or 40 = L * (L-2) * (L-2)

OpenStudy (anonymous):

I got to 40 = x^3 - 4x^2 + 4x. what i have to do next ?

OpenStudy (alekos):

we have to solve a cubic polynomial to get a value for x x^3 - 4x^2 +4x -40 = 0

OpenStudy (alekos):

By hand I worked out that the root had to be somewhere between 4 and 5, but i had to use a cubic equation calculator for this otherwise it's too cumbersome. x = 4.8669 or x = 4.87 to the nearest hundredth i'll let you work out the dimensions